1.  Cummins India Ltd has the following capital structure, which it considers optimal:

Debt 25%

Preference Shares 10%

Equity shares 65%

Total 100%

 

 Applicable tax rate for the company is 25%. Risk free rate of return is 6%, average equity market investment has expected rate of return of 12%. The company’s beta is 1.10. Following terms would apply to new securities being issued as follows:

1.  New preference can be issued at a face value of Rs. 100 per share, dividend and cost of issuance will be Rs. 10 per share and Rs. 2 per share respectively.

2.  Debt will bear an interest rate of 9%.

 

 

Calculate

  1. component cost of debt, preference shares and equity shares assuming that the company does not issue any additional equity shares.
  2. WACC.

Note - Proper calculations required in each step


1
Expert's answer
2021-05-20T13:53:03-0400

 1.

cost of debt=effective debt rate(1-tax rate)

        "=0.09(1-0.25)=0.0675=6.75\%"


cost of equity=risk free return + [beta "\times" (average market return-risk free return)]

               "=0.06+[1.10(0.12-0.06)]=0.126=12.6\%"


cost of preferred stock"=\frac{dividend}{cost \space of\space preferred\space share}"


                       "=\frac{10}{100}=0.10=10\%"



2.


"WACC=(weight\space of\space equity \times cost \space of \space equity)+(weight \space of \space debt \times cost\space of \space debt)+(weight \space of \space preferred \space stock \times cost \space of \space preferred \space stock)"


"=(0.65\times 12.6)+(0.25\times6.75)+(0.1\times10)=10.8775\%"


Hence WACC = 10.8775%